# Evaluation of machine-learning models to measure individualized treatment effects from randomized clinical trial data with time-to-event outcomes

### v1

@@ Abstract
In randomized clinical trials, regression models can be used to explore the relationships between patients’ variables (e.g., clinical, pathological or lifestyle variables, and also biomarker or genomics data) and the magnitude of treatment effect. Our aim is to evaluate the value of flexible machine learning models that can incorporate interactions and nonlinear effects of high-dimensional data to estimate individualized treatment recommendations in the setting of such trials with time-to-event outcomes. We compare survival models based on neural networks (CoxCC and CoxTime) and random survival forests (Interaction Forests). A Cox model, including an adaptive LASSO penalty, is used as a benchmark. Specific metrics for individualized treatment recommendations are used: the C-for-Benefit, the E50-for-Benefit, and RMSE for treatment benefit. We conduct an extensive simulation study using 2 different data generation processes incorporating nonlinearity and interactions up to the third order. The models are applied to gene expression and clinical data from 2 breast cancer studies. The machine learning-based methods show reasonable performances on the simulation data sets, especially in terms of discrimination for Interaction Forests and calibration for the neural networks. They can be used to evaluate individualized treatment effects from randomized trials when nonlinear and interaction effects are expected to be present.
@@ 1 Introduction
Precision medicine consists of identifying which intervention is likely to be most beneficial for a given patient. It is defined by the European Society for Medical Oncology (Yates et al., 2018) as a "healthcare approach with the primary aim of identifying which interventions are likely to be of most benefit for which patients based upon the features of the individual and their disease". It focuses on characterizing the variability in patients’ response to treatment and estimating individualized treatment benefits. In randomized clinical trials (RCTs), in which patients are randomly assigned to two treatment groups: one that receives the experimental treatment, often called the treatment group, and a second group that does not receive the treatment, often called the control group, treatment benefit is typically calculated as an average across the study population. In an RCT, an individual treatment benefit can be defined as the difference in survival outcomes for a given patient at a pre-specified time point with and without the treatment. The assessment of this treatment benefit is based on a counterfactual approach(Rubin, 1974), as both outcomes cannot be observed for a given individual.
In this work, we explore the potential of two families of machine learning methods to estimate individual patient treatment benefits from data in a high-dimensional setting and in the context of an RCT. We apply $2$ strategies for the feedforward neural networks (FNNs) based on a specific loss function in a continuous time framework (CoxCC and CoxTime) (Kvamme et al., 2019), and Interaction Forests (IF), an algorithm based on diversity forests that accounts for bidirectional interactions. We benchmark these methods with a penalized Cox Proportional Hazards (CoxPH) model (Cox, 1972) with an adaptive LASSO penalty, as previously recommended by Ternès et al. (2017).
Our study examines the potential of the methods to estimate individual patient treatment benefit from data in a high-dimensional setting, when non-linear and interaction effects are present. We apply calibration and discrimination measures specifically designed for the potential outcome framework and adapt some of these measures to a time-to-event outcome.
The performances of the different models are evaluated and compared using a simulation study with two different data generation processes. To overcome the linear framework without interactions, we simulate data with nonlinearity and interactions up to the third order. We also present the results on two breast cancer patient data sets with high-dimensional gene expression, evaluating the effect of adding either taxane chemotherapy or trastuzumab to a standard chemotherapy treatment.
@@ 2 Related work
In an RCT, the analysis focuses on the treatment effect, which refers to the impact of a treatment on a specific outcome variable. This variable can be of various types, such as a binary outcome or time-to-event. Different methods have been suggested to propose a measure of the treatment effect. With the FNN model DeepSurv, Katzman et al. (2018) provide a personalized recommendation method derived from randomized data, which may predict survival in a subgroup of patients. Klaveren et al. (2017) focus on outcome risk and define treatment benefit as the difference between outcome risk with and without therapy. They define treatment benefit in $2$ different ways. Observed treatment benefit is the difference in survival outcome between $2$ patients with the same predicted benefit but with different treatment assignments. The predicted treatment benefit is the predicted risk with a treatment for a given patient minus the predicted risk with an alternative treatment option for the same patient.
In the counterfactual framework, a matching procedure should be implemented to compute observed treatment benefit. Different methods can create pairs of patients with similar profiles but different treatment assignments. Patients can be matched on predicted benefit (Klaveren et al., 2017): $2$ patients are paired if they have close predicted benefit and discordant treatment assignments. Klaveren et al. (2017) compare this procedure with an alternative matching based on covariates. This second matching procedure is also applied by Maas et al. (2023) They use the Mahalanobis distance between patients’ characteristics to create pairs of patients.
Klaveren et al. (2017) study binary outcomes and predicted risk, discussing how to extend their measure to time-to-event data. Rekkas et al. (2023) estimate individualized treatment effects from an RCT, comparing various risk-based approaches with a binary outcome. Bouvier et al. (2024) investigate methods for evaluating individualized treatment effects with time-to-event outcomes using individual participant data meta-analysis, highlighting the potential of meta-analysis in predicting individualized treatment outcomes.
@@ 3.1 Notations
Let the random variable $T\in\mathbb{R}^{+}$ represent the survival time, that is the time between the starting point and the occurrence of a given event (e.g. the time between a patient’s treatment assignment and death). The survival function at time $t$, $S(t)$, is defined as:
@@
[…]
Often, $T$ is not observed for all individuals: time-to-event data is censored. Survival data is said to be right-censored if the individual leaves the study before the end point or if the event did not happen before the end of the follow-up. In the case of right-censoring, we define $(C_{i})_{i=1,\dots,n}$ the independent and identically distributed censoring times of $n$ individuals. We can set $\{\tilde{T}_{i},X_{i},D_{i})\}_{i=1,\dots,n}$ where $\tilde{T}=\min(T_{i},C_{i})$ is the time until death or censoring, $X_{i}=(X_{i,1},\dots,X_{i,p})^{T}\in\mathbb{R}^{p}$ denotes a $p$-dimensional vector of variables, and $D_{i}=\mathbb{I}\{\tilde{T}_{i}=T_{i}\}$ is the censoring indicator. We suppose that the survival time $T_{i}$ is independent of the censoring time conditionally on the variables vector $X_{i}$ for $i=1,\dots,n$. $G(t)$ can be defined as the estimate of the censoring survival function:
@@ 3.2 Models
In this paper, we compare $3$ survival-adapted machine learning algorithms and investigate their ability to predict treatment effects. More specifically, we compare $2$ FNNs that use specific loss functions to handle censored observations with IF, a specific type of random survival forest (RSF) that models interaction effects. These models are benchmarked with a penalized CoxPH with linear effects and the LASSO penalty.
@@ 3.2.1 Benchmark method : CoxPH with adaptive LASSO penalty
With high dimensional data, the model can become non-identifiable. A LASSO penalty can be introduced to select only the variables with the strongest effects on the outcome of interest. First introduced in the context of linear regression by Tibshirani (1996) and then adapted to survival analysis (Tibshirani, 1997), the LASSO penalty corresponds to the $L_{1}$ norm of the regression coefficients. The penalization term is denoted by:
$pen(\lambda)=\lambda||\beta||_{1}=\lambda\left(\sum_{j=1}^{p}|\beta_{j}|\right).$
The model’s degree of parsimony or complexity depends on the penalization parameter $\lambda$: it varies from 00 (complete model including all biomarkers) to +$\infty$+ ∞ (null model with no biomarkers). $\lambda$ is often estimated using cross-validation (CV).
In certain scenarios, LASSO can be inconsistent for variable selection. Here, we implement adaptive lasso (ALASSO) (Fan and Lv, 2008), a method based on an additive penalization term for a more conservative choice of the shrinkage parameter lambda. With ALASSO, adaptive weights are used for penalizing difference coefficients in the $L_{1}$ penalty. The penalization term becomes:
@@
where $w_{j}=\frac{1}{|\hat{\beta}_{j}|}$ are biomarker-specific weights, and $\hat{\beta}_{j},j=1,\dots,p$ are are estimated by fitting a preliminary regular LASSO. The non-negative penalty parameter $\lambda$ for the adaptive lasso is chosen using the maximum cross-validated log-likelihood method.
@@ 3.2.2 Machine learning algorithms adapted to time-to-event outcomes
The hypotheses of the linear CoxPH model are restrictive, for instance, when nonlinear and complex interactions exist in the data. Here, we focus on artificial neural networks and random forests, as these algorithms can model complex patterns.
Kvamme et al. (2019) introduced a method called CoxCC that uses a loss function based on a case-control approximation. They proposed to randomly sample a new set of controls at each iteration instead of keeping control samples fixed. The loss is written as:
@@
with ~$\tilde{\mathcal{R}}_{i}$ a subset of the risk set $\mathcal{\mathcal{R}}_{i}$ at time $t$ including individual $i$. $i$ represents the case, and the $j$’s are the controls sampled from the risk set. $\phi$ represents the transformation operated by the FNN on the input variables. CoxCC can be trained with this specific loss (Equation 3) using a mini-batch gradient descent algorithm.
@@
The FNNs are compared with IF (Hornung and Boulesteix, 2022), a type of random survival forest specifically designed to model quantitative and qualitative interaction effects in bivariate splits. $2$ variables are said to interact if the effect of one variable on the outcome depends on the value of the other variable. Here, we focus on biomarker-by-treatment interactions in the setting of an RCT. Two types of interactions are accounted for: the quantitative interactions and the qualitative interactions, as categorized by Peto (1982). For a quantitative interaction, the treatment effect differs quantitatively depending on the biomarker value, but it remains in the same direction conditional on biomarker values (e.g., the treatment effect increases when the biomarker value increases). Conversely, when considering a qualitative interaction, the direction of the treatment effect depends on the biomarker value. As they are based on multivariable splitting, IFs are specifically developed to model these interactions.
@@ 3.2.3 Hyperparameter search
For the FNN models, we perform the hyperparameter search in the context of a 5-fold CV applied on the training set. First, the data is split into a training and a test set. Then the 5-fold CV is applied to the training set. It consists in randomly splitting the training data into $K=5$ folds. The model is trained using $K-1$ folds and validated on the remaining fold, outputting a score. The training is repeated until each fold is used as a validation set. The average of the recorded scores is the performance metric of the model. Finally, the entire training set is used to train each network with the previously selected set of hyperparameters. The results are outputted on the test set.
We use the Tree-Parzen algorithm (Bergstra et al., 2011) to select hyperparameters iteratively in an informed manner. We define a search space for each hyperparameter with specific distribution and boundary values. These values are described in Table 7. A set of hyperparameters is randomly sampled, and the model is scored on each of the five validation folds. These five validation scores are averaged, and a new set of hyperparameters is sampled based on the value of the average score. The sampling of hyperparameters sets is repeated $200$ times.
For the application to the two patient data sets, we perform a double 5-fold CV on the entire dataset, because we have relatively small datasets and want to mimic an external test set. This strategy was first illustrated by Matsui et al. (2012). First, the real patient cohort is split into five folds: this is the outer loop. Then, we select one of the five folds as a test set and perform a 5-fold CV on the remaining data for each hyperparameter set: this is the inner loop. We choose the hyperparameters configuration with the minimum average validation loss obtained on the five folds of the inner loop. Finally, we fit the model with these optimal hyperparameters on the four folds of the outer loop and calculate predictions on the remaining test fold of the outer loop. This procedure is repeated on all the folds of the outer loop.
@@ 3.3 Data
The models are trained on 2 types of data: simulated data and real patient data sets. The synthetic data are generated to examine the operating characteristics and performances of the modeling techniques, while 2 high-dimensional patient clinical trial data sets are used for illustrative purposes.
@@ 3.3.1 Simulation study
Rothwell (2005) put forward that the only reliable approach for assessing the predictiveness of biomarkers is to test their interaction with the treatment and introduce a full biomarker-by-treatment interaction model. We can simulate data based on this model:
@@
In this model, we add non-linearity with the choice of the function $f$, which enables the building of complex interactions with the treatment. Our model extends the model from Haller et al. (2019) which considered the case of a single biomarker.
Here, $\alpha_{\tau}$ is chosen as $\alpha_{\tau}=ln(0.75)\approx-0.288$. We set $\beta_{i1}=0.288$, $f_{i}(X)=(2X-1)^{2}$ and $\beta_{i2}=-0.9$. The biomarkers are generated from a uniform distribution: $X_{i}\sim\mathcal{U}_{[0,1]}$. Since we intend to simulate an RCT, a total of $n$ patients per data set are randomly assigned with a 1:1 ratio to the experimental or control arm, with $P(\tau=\frac{1}{2})=P(\tau=-\frac{1}{2})=\frac{1}{2}$.
@@
A cohort of size $n$ individuals ($n\in\{2,000;20,000\}$) is generated for each set, split into $50\%$ for the training set and $50\%$ for the test set. We generate independent censoring, and consider two censoring rates: a moderate censoring rate of $20\%$ and a large censoring rate of $50\%$. $q=20$ variables are generated for each set. Among $q$, only $p=10$ biomarkers are really prognostic. For each scenario, $100$ replications are performed.
In the second data generation process, nonlinear interactions are simulated based on a Friedman’s random function generator (Friedman, 2001) in an accelerated failure time (AFT) model. Friedman’s random function generator allows us to generate random functions with second-order interactions between the survival data and the explanatory variables and strong nonlinear effects. It is based on an AFT model, that assumes that the relationship between the logarithm of survival time $T$ and the variables is linear. The random function generator is defined as:
@@
Henderson et al. (2020) adapt the flexible random function generator to the analysis of treatment effect by introducing a second generator interacting with the treatment variable. For this purpose, the AFT-Friedman model can be rewritten as:
@@
$S(t|X)=\exp\Bigg{(}-\int_{0}^{t}h(s|X)ds\Bigg{)}=\exp\Bigg{(}-H_{0}(t\exp(m_{1}(X)+\tau\times m_{2}(X))\Bigg{)}$ (10)
A cohort of size $n$ individuals ($n\in\{2,000;20,000\}$) is generated for each set, split into $50\%$ for the training set and $50\%$ for the test set. We generate independent censoring, and considered a moderate censoring rate of $20\%$. For each scenario, $100$ replications are performed.
3.4 Application on cancer data sets
We apply the methods to gene expression data from a meta-analysis of neoadjuvant trials that included $614$ breast cancer patients treated by anthracyclines alone or anthracyclines plus taxanes (Ternès et al., 2018). We also use data from an RCT including $1,574$ patients, which evaluates the effect of adjuvant trastuzumab in breast cancer (Pogue-Geile et al., 2013).

### v2

@@ Abstract
Objective: In randomized clinical trials, prediction models can be used to explore the relationships between patients’ variables (e.g., clinical, pathological, or lifestyle variables, and also biomarker or genomic data) and treatment effect magnitude. Our aim was to evaluate flexible machine learning models capable of incorporating interactions and nonlinear effects from high-dimensional data to estimate individualized treatment recommendations in trials with time-to-event outcomes.
Methods: We compared survival models based on neural networks (CoxCC and CoxTime) and random survival forests (Interaction Forests) against a Cox proportional hazards model with an adaptive LASSO (ALASSO) penalty as a benchmark. For individualized treatment recommendations in the survival setting, we adapted metrics originally designed for binary outcomes to accommodate time-to-event data with censoring. These adapted metrics included the C-for-Benefit, the E50-for-Benefit, and the root mean squared error for treatment benefit. An extensive simulation study was conducted using two different data generation processes incorporating nonlinearity and interactions. The models were applied to gene expression and clinical data from three cancer clinical trial data sets.
Results: In the first data generation process, neural networks outperformed ALASSO in terms of calibration while the Interaction Forests showed superior C-for-benefit performance. In the second data generation process, both machine learning methods outperformed the benchmark linear ALASSO method across discrimination, calibration, and RMSE metrics. In the cancer trial data sets, the machine learning methods often performed better than ALASSO, particularly IF in terms of C-for-benefit, and either a neural network or IF for calibration measures addressing treatment benefit. Conclusion: Machine learning-based survival models can efficiently estimate conditional average treatment effects from randomized trials when strong nonlinear and interaction effects are expected. These flexible approaches offer valuable alternatives to traditional regression methods for developing treatment recommendation rules from clinical trials with time-to-event outcomes and medium to high-dimensional data set sizes.
@@ 1 Introduction
Precision medicine involves identifying the intervention that is most likely to be beneficial for a given patient. It is defined by the European Society for Medical Oncology (Yates et al., 2018) as a "healthcare approach with the primary aim of identifying which interventions are likely to be of most benefit for which patients based upon the features of the individual and their disease". It focuses on characterizing the variability in patients’ response to treatment and estimating treatment benefits.
In randomized controlled trials (RCTs), in which patients are randomly assigned to two treatment groups: one that receives the experimental treatment, often called the treatment group, and a second group that does not receive the treatment, often called the control group, the treatment benefit is typically calculated as an average across the study population. In an RCT, an individual treatment benefit can be defined as the difference in survival outcomes for a given patient at a pre-specified time point with and without the treatment. The assessment of this treatment benefit is based on a counterfactual approach (Rubin, 1974), as both outcomes cannot be observed for a given individual.
In this work, we explore the potential of two families of machine learning methods to estimate conditional average treatment effects from data in a medium to high-dimensional setting and in the context of an RCT, aiming to predict invidualized treatment effects. We apply $2$ strategies for the feedforward neural networks (FNNs) based on a specific loss function in a continuous time framework (CoxCC and CoxTime) (Kvamme et al., 2019), and Interaction Forests (IF), an algorithm based on diversity forests that accounts for bidirectional interactions. We benchmark these methods with a penalized Cox Proportional Hazards (CoxPH) model (Cox, 1972) with an adaptive LASSO penalty, as previously recommended by Ternès et al. (2017) for this setting.
Our study examines the potential of these methods to estimate individual patient treatment benefits from data in a high-dimensional setting, where nonlinear and interaction effects are present. We apply calibration and discrimination measures specifically designed for the potential outcome framework and adapt some of these measures to a time-to-event outcome.
The performances of the different models are evaluated and compared using a simulation study with two different data generation processes, which simulate data with the presence of nonlinearity and interactions. We also present the results of two case studies from clinical trials in cancer utilizing gene expression data.
@@ 2 Related work
In an RCT, the main analysis focuses on the average treatment effect, which refers to the impact of a treatment on a specific outcome variable. This outcome can be of various types, such as a binary outcome or time-to-event. In this work, we are interested in the heterogeneity of the treatment effect based on patient characteristics. An overview of recent developments in statistical methods for assessing the heterogeneity of treatment effects, with a particular focus on subgroup identification and the estimation of individualized treatment strategies is provided in Lipkovich et al. (2024). For survival outcomes, the FNN model DeepSurv, Katzman et al. (2018) provide a personalized recommendation method derived from randomized data, which may predict survival in a subgroup of patients. Klaveren et al. (2017) focus on outcome risk and define treatment benefit as the difference between outcome risk with and without therapy. They define treatment benefit in $2$ different ways. Observed treatment benefit is the difference in survival outcome between $2$ patients with the same predicted benefit but with different treatment assignments. The predicted treatment benefit is the predicted risk with a treatment for a given patient minus the predicted risk with an alternative treatment option for the same patient.
In the counterfactual framework, a matching procedure should be implemented to compute the observed treatment benefit. Different methods can create pairs of patients with similar profiles but different treatment assignments. Patients can be matched on predicted benefit (Klaveren et al., 2017): $2$ patients are paired if they have close predicted benefit and discordant treatment assignments. Klaveren et al. (2017) compare this procedure with an alternative matching based on covariates. This second matching procedure is also applied by Maas et al. (2023). They use the Mahalanobis distance between patients’ characteristics to create pairs of patients.
Klaveren et al. (2017) study binary outcomes and predicted risk, discussing how to extend their measure to time-to-event data. Rekkas et al. (2023) estimate individualized treatment effects from an RCT, comparing various risk-based approaches with a binary outcome. Bouvier et al. (2024) investigates methods for evaluating individualized treatment effects with time-to-event outcomes using individual participant data meta-analysis, highlighting the potential of multiple clinical trial data sets in a meta-analysis to predict individualized treatment effects.
@@ 3.1 Notations
Let the random variable $T\in\mathbb{R}^{+}$ represent the survival time, that is, the time between the starting point and the occurrence of a given event (e.g., the time between a patient’s treatment assignment and death). The survival function at time $t$, $S(t)$, is defined as:
@@
with $H(t)=\int_{0}^{t}h(s)ds$ , and $h(t)$ the hazard rate. $H_{0}$ is the cumulative hazard at baseline for time $t=0$.
Often, $T$ is not observed for all individuals: time-to-event data is censored. Survival data are said to be right-censored at the last date of follow-up, meaning that no event has occurred. In the case of right-censoring, we define $(C_{i})_{i=1,\dots,n}$ the independent and identically distributed censoring times of $n$ individuals. We can set $\{\tilde{T}_{i},X_{i},D_{i}\}_{i=1,\dots,n}$ where $\tilde{T}=\min(T_{i},C_{i})$ is the time until death or censoring, $X_{i}=(X_{i,1},\dots,X_{i,p})^{T}\in\mathbb{R}^{p}$ denotes a $p$-dimensional vector of variables, and $D_{i}=\mathbb{I}\{\tilde{T}_{i}=T_{i}\}$ is the censoring indicator. We suppose that the survival time $T_{i}$ is independent of the censoring time conditionally on the vector of variables $X_{i}$ for $i=1,\dots,n$. $G(t)$ can be defined as the estimate of the censoring survival function:
@@ 3.2 Models
In this paper, we compare $3$ survival-adapted machine learning algorithms and investigate their ability to estimate conditional average treatment effects. The treatment effect can be defined as the effect of treatment $\tau$ on the outcome variable $Y$. Since $Y$ is the observed outcome, we have:
In an RCT, assuming that $Y(1)$ and $Y(0)$ are independent from $\tau|X$, the conditional average treatment effect can be defined as:
with $X$ the characteristics of the patient.
More specifically, we compare $2$ FNNs that use specific loss functions to handle censored observations with IF, a specific type of random survival forest (RSF) that models interaction effects. These models are benchmarked with a penalized CoxPH with linear effects and the LASSO penalty.
@@ 3.2.1 Benchmark method: CoxPH with adaptive LASSO penalty
With high-dimensional data, the model can become non-identifiable. A LASSO penalty can be introduced to select only the variables with the strongest effects on the outcome of interest. First introduced in the context of linear regression by Tibshirani (1996) and then adapted to survival analysis (Tibshirani, 1997), the LASSO penalty corresponds to the $L_{1}$ norm of the regression coefficients. The penalization term is denoted by:
[…]
The model’s degree of parsimony or complexity depends on the penalization parameter $\lambda$: it varies from 0 (complete model including all biomarkers) to $\infty$ (null model with no biomarkers). $\lambda$ is often estimated using cross-validation (CV).
In certain scenarios, LASSO can be inconsistent for variable selection. Here, we implement adaptive lasso (ALASSO) (Fan and Lv, 2008), a method based on an additive penalization term, which provides a more conservative choice of the shrinkage parameter $\lambda$. With ALASSO, adaptive weights are used for penalizing difference coefficients in the $L_{1}$ penalty. The penalization term becomes:
@@
where $w_{j}=\frac{1}{|\hat{\beta}_{j}|}$ are biomarker-specific weights, and $\hat{\beta}_{j},j=1,\dots,p$ are estimated by fitting a preliminary regular LASSO. The non-negative penalty parameter $\lambda$ for the adaptive lasso is chosen using the maximum cross-validated log-likelihood method.
@@ 3.2.2 Machine learning algorithms adapted to time-to-event outcomes
The hypotheses of the linear CoxPH model are restrictive, for instance, when nonlinear and complex interactions exist in the data. Here, we focus on artificial neural networks and random forests, as these algorithms are well-suited to model complex interaction patterns.
Kvamme et al. (2019) introduced a method called CoxCC that uses a loss function based on a case-control approximation. They proposed randomly sampling a new set of controls at each iteration instead of keeping the control samples fixed. The loss is written as:
@@
with $\tilde{\mathcal{R}}_{i}$ a subset of the risk set $\mathcal{\mathcal{R}}_{i}$ at time $t$ including individual $i$. $i$ represents the case, and the $j$’s are the controls sampled from the risk set. $\phi$ represents the transformation operated by the FNN on the input variables. CoxCC can be trained with this specific loss (Equation 3) using a mini-batch gradient descent algorithm.
@@
In this paper, we compare the FNNs with IF (Hornung and Boulesteix, 2022), a type of random survival forest specifically designed to model quantitative and qualitative interaction effects in bivariate splits. $2$ variables are said to interact if the effect of one variable on the outcome depends on the value of the other variable. Here, we focus on biomarker-by-treatment interactions in the setting of an RCT. Two types of interactions are accounted for, quantitative and qualitative interactions, as categorized by Peto (1982). For a quantitative interaction, the treatment effect varies quantitatively depending on the biomarker value, but remains in the same direction, conditional on the biomarker value (e.g., the treatment effect increases as the biomarker value increases). Conversely, when considering a qualitative interaction, the direction of the treatment effect depends on the value of the biomarker.
@@ 3.2.3 Hyperparameter search
For the FNN models, we perform the hyperparameter search in the context of a 5-fold CV applied to the training set. Then the 5-fold CV is applied to the training set. It consists of randomly splitting the training data into $K=5$ folds. The model is trained using K1K-1 folds and validated on the remaining fold. The training is repeated until each fold is used as a validation set. Finally, the entire training set is used to train each network with the previously selected set of hyperparameters. We use the Tree-Parzen algorithm (Bergstra et al., 2011) to select hyperparameters iteratively in an informed manner. We define a search space for each hyperparameter with specific distribution and boundary values. These values are described in Table 6. A set of hyperparameters is randomly sampled, and the model is scored on each of the five validation folds. These five validation scores are averaged, and a new set of hyperparameters is sampled based on the value of the average score. The sampling of hyperparameter sets is repeated $200$ times.
@@
For the three case studies, we perform a double 5-fold CV on the entire dataset, because we have relatively small datasets and want to mimic an external test set. This strategy was first illustrated by Matsui et al. (2012). First, the real patient cohort is split into five folds, which constitute the outer loop. Then, we select one of the five folds as a test set and perform a 5-fold CV on the remaining data for each hyperparameter set: this is the inner loop. We choose the hyperparameter configuration with the minimum average validation loss obtained on the five folds of the inner loop. Finally, we fit the model with these optimal hyperparameters on the four folds of the outer loop and calculate predictions on the remaining test fold of the outer loop. This procedure is repeated on all the folds of the outer loop.
@@ 3.3 Data
The models are trained on two types of data: simulated data and real patient data sets. The synthetic data are generated to examine the operating characteristics and performances of the modeling techniques, while two high-dimensional patient clinical trial data sets are used for illustrative purposes.
@@ 3.3.1 Simulation study
Rothwell (2005) put forward that the only reliable approach for assessing the predictiveness of biomarkers is to test their interaction with the treatment and introduce a full biomarker-by-treatment interaction model. Let $\tau$ denote the treatment indicator, with $\tau=1$ for patients receiving the treatment and $\tau=0$ for those in the control group, and let $X=(X_{1},\dots,X_{p})^{T}\in\mathbb{R}^{p}$ denote the vector of patient covariates. We can simulate data based on this model:
@@
In this model, we introduce nonlinearity through the choice of the function $f$, which enables the construction of complex interactions with the treatment. Our model extends the model from Haller et al. (2019), which considered the case of a single biomarker.
Here, $\alpha_{\tau}$ is chosen as $\alpha_{\tau}=ln(0.75)\approx-0.288$. We set $\beta_{i1}=0.288$, $f_{i}(X)=(2X-1)^{2}$ and $\beta_{i2}=-0.9$. The biomarkers are generated from a uniform distribution: $X_{i}\sim\mathcal{U}_{[0,1]}$. Since we intend to simulate an RCT, a total of $n$ patients per data set are randomly assigned with a 1:1 ratio to the experimental or control arm, with $P(\tau=1)=P(\tau=0)=\frac{1}{2}$.
@@
A cohort of $n$ individuals ($n\in\{2,000;20,000\}$) is generated for each set, split into a $50\%$ training set and a $50\%$ test set. We generate independent censoring, and consider two censoring rates: a moderate censoring rate of $20\%$ and a high censoring rate of $50\%$. $q=20$ variables are generated for each set. Among the $q$ biomarkers, only $p=10$ are truly prognostic. For each scenario, $100$ replications are performed.
In the second data generation process, nonlinear interactions are simulated using Friedman’s random function generator (Friedman, 2001) within an accelerated failure time (AFT) model. Friedman’s random function generator enables us to generate random functions with second-order interactions between the survival data and the explanatory variables, as well as strong nonlinear effects. It is based on an AFT model that assumes a linear relationship between the logarithm of survival time $T$ and the variables. The random function generator is defined as:
@@
Henderson et al. (2020) adapt the flexible random function generator specifically to treatment effects by introducing a second generator interacting with the treatment variable. For this purpose, the AFT-Friedman model can be rewritten as:
@@
$S(t|X)$ $=\exp\Bigg(-\int_{0}^{t}h(s|X)ds\Bigg)$
$=\exp\Bigg(-H_{0}(t\exp(m_{1}(X)+\tau\times m_{2}(X))\Bigg)$ (10)
A cohort of $n$ individuals ($n\in\{2,000;20,000\}$) is generated for each set, split into $50\%$ for the training set and $50\%$ for the test set. We generate independent censoring, and consider a moderate censoring rate of $20\%$. For each scenario, $100$ replications are performed.
3.4 Cases studies: application on cancer data sets
We apply the methods to gene expression data from a meta-analysis of neoadjuvant trials that included $614$ breast cancer patients treated by anthracyclines alone or anthracyclines plus taxanes (Ternès et al., 2018). We also use gene expression data from an RCT including $1,574$ patients, which evaluated the effect of adjuvant trastuzumab in early breast cancer (Pogue-Geile et al., 2013), and finally gene signature data an RCT in renal cancer with $n=886$ patients, which assessed first-line avelumab plus axitinib versus sunitinib in previously untreated patients with advanced renal cell carcinoma (Motzer et al., 2020).